OpenStudy (anonymous):

A student at a university makes money by buying and selling used cars. Charles bought a used car and later sold it for a 20% profit. If he sold it for $4590, how much did Charles pay for the car?

OpenStudy (anonymous):

cost price = 100/(100+profit%) * S.P so CP=100/120 *4590 CP=3825

OpenStudy (perl):



OpenStudy (perl):